# Gradual smoothing: strong hypercontractivity and logarithmic Sobolev inequalities

### v1

@@ 1.4. Outline of the paper
In Section 2 we introduce several necessary concepts and recall some useful tools already available in the literature. Section 3 is devoted to proving the equivalence between smoothing effects and families of logarithmic Sobolev inequalities. In Section 4 we focus on the model operator $\log(I-\Delta)$, to which we apply the previous equivalence result in order to obtain a new family of logarithmic Sobolev inequalities.  Section 5 addresses general operators, making extensive use of the results from the preceding sections. In Section 6 we prove eventual ultracontractivity under the additional hypothesis of translation invariance. Finally, Section 7 contains additional comments and possible extensions. For completeness, an Appendix summarizes the relevant existence theory.
@@ 3. Hypercontractivity and logarithmic Sobolev inequalities
There is a reason to talk about pp–hypercontractivity, and not about pp–supercontractivity or pp–ultracontractivity. In the usual instances found in the literature, two pp–hypercontractivities with two different values of $p$ imply each other quantitatively, since they are both equivalent to the standard 22–logarithmic Sobolev inequality. However, to show strong hypercontractivity we need a family of independent pp–logarithmic Sobolev inequalities; therefore, the pp–hypercontractivities are also expected to be quantitatively independent; see the discussion below. The case of supercontractivity and ultracontractivity is different, since through an interpolation and iteration argument in (3.2) and (3.3), standard in the literature, one could substitute the 22–norm by any pp–norm.
@@
Let now $u$ be a solution to ($\textup{P}_{\mathcal{L}}$). Recall that we may assume that $u\geq 0$, and we will do so in the sequel. Also, consider all the integrals to be over $\mathbb{R}^{N}$. Differentiating $\int u^{q(t)}(t)$, we formally obtain
@@
An approximation argument using Section 3 shows that (3.22) is also true in $L^{1}$, hence (3.21) holds in $L^{1}$. Finally, we integrate to get (3.15). ∎
@@
Now, let us prove a reciprocal result, which requires stronger hypotheses. While we required the pp–hypercontractivity to hold in order to prove the pp–logarithmic Sobolev inequality, we need all the rr–logarithmic Sobolev inequalities for $r\geq p$ to prove the pp–hypercontractivity.
@@
Fix $p\geq 2$. Assume that the rr–logarithmic Sobolev inequalities (3.10) hold for all $r\geq p$. Then $\mathcal{L}$ is hypercontractive with the family of exponents satisfying
@@
(3.37) $U^{\prime}(t)+s(t)\int u^{s(t)-1}\mathcal{L}u\leq\frac{s^{\prime}(t)}{s(t)}\left\{C(s(t))U(t)+D(s(t))\mathcal{E}\big(u^{s(t)-1}(t),u(t)\big)+U(t)\log\left(U(t)\right)\right\}.$
@@
With the two previous results we notice that, given $p\geq 2$, satisfying the rr–logarithmic Sobolev inequalities for every $r\geq p$ is equivalent to being rr–hypercontractive for every $r\geq p$. We would like to extend the result to $p\in(1,2)$. This will be done through a duality argument, which requires to show that the semigroup $e^{-t\mathcal{L}}$ is self-adjoint, an immediate consequence of $\mathcal{L}$ being self-adjoint.
@@
Let $u$ and $v$ be solutions to problem ($\textup{P}_{\mathcal{L}}$) with initial data $u_{0},v_{0}\in L^{2}(\mathbb{R}^{N})$ respectively. We want to prove that
@@
Hence, we have that $t_{q^{\prime}}^{p^{\prime}}=t_{p}^{q}$, and we can define $A_{q^{\prime}}(t)=A_{p}(t)$. Therefore $\mathcal{L}$ is $p$-hypercontractive for every $1<p\leq 2$.
@@
The function $q$ solution of the ODE satisfies the conditions of Theorem 3.11. Condition (3.51) is nothing but Osgood’s blow-up condition for the Cauchy problem associated to $q$, written in terms of $D$ instead of $h$. ∎
@@
Under the hypotheses of Theorem 3.12, the operator $\mathcal{L}$ is eventually ultracontractive if
@@
(3.56) $\|u(t)\|_{q(t)}=A_{p}(t)\|u_{0}\|_{p},\quad\text{where }A_{p}(t)=\frac{q(t)}{p}.$
Hence, $\lim_{t\to t_{\infty}}A_{p}(t)=\infty$. This means that even though $q(t)$ is exploding to infinity at $t_{\infty}$, we cannot show that $u$ becomes bounded through this method. In fact, along the paper we will find that solutions in general will not become bounded at $t=t_{\infty}$. Nevertheless, they will become bounded for $t>t_{\infty}$, at least in the translation-invariant setting.
@@ 4. The $\log(I-\Delta)$ operator
To apply the results of the previous section, we will work with near-zero order operators, also known as geometrically stable operators. The idea is to work with an operator $\mathcal{L}$ with symbol $m(\xi)\sim\log(\xi)$ for $\xi\sim\infty$, since these are the kind of operators expected to have hypercontractivity, but not ultracontractivity; see Section 6. If we use $m(\xi)=\log(1+|\xi|^{2})$, we find the model operator for this work, $\mathcal{L}=\log(I-\Delta)$. This operator has the advantage of being a subordinated operator to the Laplacian. It is defined as
@@
It can also be understood, though we do not use it, as the derivative at $\alpha=0$ of the Bessel operators $(I-\Delta)^{\alpha}$, i.e.,
@@ 4.1. Representation in terms of a Lévy kernel
On the other hand, using Laplace’s method, see for instance [38], we can calculate, for large $|z|$,
$\begin{array}[]{rll}J(z)&=\displaystyle(4\pi)^{-\frac{N}{2}}\int_{0}^{\infty}t^{-\frac{N}{2}-1}e^{-\frac{|z|^{2}}{4t}-t}\,dt=\int_{0}^{\infty}g_{1}(t)e^{-g_{2}(z,t)}\,dt\\[8.53581pt] &\displaystyle=g_{1}(t_{0})e^{-g_{2}(z,t_{0})}\left(\int_{0}^{\infty}e^{-\frac{1}{2}\frac{\partial^{2}g_{2}}{\partial t^{2}}(z,t_{0})s^{2}}\,ds+o(1)\right)\\[8.53581pt] &\displaystyle=g_{1}(t_{0})e^{-g_{2}(z,t_{0})}\left(\left(\frac{2\pi}{\frac{\partial^{2}g_{2}}{\partial t^{2}}(z,t_{0})}\right)^{1/2}+o(1)\right),\end{array}$~
@@
(4.8) $|z|^{\frac{N+1}{2}}e^{|z|}J(z)=\left((2\pi)^{-\frac{N-1}{2}}+o(1)\right)\quad\text{as }|z|\to\infty.$
The behaviour $J(z)\sim|z|^{-N}$ for $|z|\sim 0$ is what makes us call this operator a near-zero order operator. This fact is of great importance, since $\log(I-\Delta)$ will be the model for all operators whose kernel behaves that way. The behaviour at infinity is also studied in [49]. In particular the integral is a representation of a Bessel function of second order.
@@ 4.3. Smoothing properties
From this it is easy to observe that $H_{t}\in L^{\frac{N}{N-2t},\infty}(\mathbb{R}^{N})$, since $\||x|^{-N+2t}\|_{\frac{N}{N-2t},\infty}=\left(\frac{\omega_{N-1}}{N}\right)^{\frac{N-2t}{N}}$. Notice that with this information we may deduce the Sobolev embeddings found in Section 4.3. We may also deduce a similar result for $L^{1}$ initial data.
@@
Even though (4.22) was useful for this qualitative result, there are hidden factors depending in $t$ in its last term. Thus, the above formula is not enough to get precise quantitative estimates. Instead, we use a simpler but, for our purposes, much more powerful approach. In (4.15) we perform the change of variables $\tau=\frac{|x|^{2}}{4s}$ , which yields
@@ 4.4. Quantitative hypercontractivity
(4.37) $\|u(t)\|_{\frac{Np}{N-2pt}}\leq A_{p}(t)\|u_{0}\|_{p}\quad\text{for }0\leq t<\frac{N}{2p}.$
@@
where $C_{N,p}(t)$ is the best possible constant for the inequality. We need the explicit expression of $C_{N,p}(t)$, or at least a sufficiently good upper bound. The best constant in weak Young’s inequality is exactly the best constant $c_{N,p,\lambda}$ for Hardy-Littlewood-Sobolev’s inequality,
@@ 4.5. Logarithmic Sobolev inequalities
We have now all the ingredients needed to establish the logarithmic Sobolev inequalities for $\log(I-\Delta)$.
@@
(4.45) $A_{2}^{\prime}(t)=A_{2}(t)\Big[-\psi\!\left(\frac{N}{2}-2t\right)-\psi\!\left(\frac{N}{2}+2t\right)-\ln(4\pi)+\frac{2}{N}\ln\!\left(\frac{\Gamma(N)}{\Gamma(\tfrac{N}{2})}\right)\Big].$
@@ 4.6. Comparison with the operator $\log(-\Delta)$
with the same function coefficient $A_{p}$ obtained in Theorem 4.7. Hence, one may try to apply Theorem 3.11 in order to obtain the family of logarithmic Sobolev inequalities for $\log(-\Delta)$. However, the expression (3.15) is not well defined for functions of the form $u(t)=R_{t}*u_{0}$. Consequently, obtaining logarithmic Sobolev inequalities —at least by our method— is not possible. In fact, $u(t)$ cannot be considered a solution in the sense of Section 2.2, since $u(t)$ may not belong to $L^{2}$ for $t>0$. It is clear that Hille-Yosida’s Theorem, Theorem 7.1, cannot be applied here, since $\log(-\Delta)$ is not a monotone operator.
Even more important, solutions $u(t)=R_{t}*u_{0}$ for $u_{0}\in L^{p}(\mathbb{R}^{N})$ stop being well-defined for $t>\frac{N}{2p}$. So the main differences between problem ($\textup{P}_{\mathcal{L}}$) for $\log(-\Delta)$ and $\log(I-\Delta)$ are:
@@ 6. Translation-invariant operators
Let $1\leq p<\infty$. If $e^{-m}\in L^{p}(\mathbb{R}^{N})$, then $\mathcal{H}_{t}\in L^{1}(\mathbb{R}^{N})\cap L^{\frac{p}{p-t}}(\mathbb{R}^{N})$ for $t\geq p/2$. In particular, $\mathcal{H}_{t}\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ for $t\geq p$.
@@
Define $\tau_{z}f(x)=f(z-x)$. Then, condition (6.5) can be rewritten as
(6.6) $|\langle H_{t_{0}},\tau_{x}f\rangle|\leq C\|\tau_{x}f\|_{p}$
for every $f\in L^{p}(\mathbb{R}^{N})$, since $\tau_{x}$ is an isometry in $L^{p}(\mathbb{R}^{N})$. Hence, renaming $g=\tau_{x}f$, and taking the supremum of $g\in L^{p}(\mathbb{R}^{N})$ such that $\|g\|_{p}=1$, we conclude what we wanted. ∎
If $m$ grows logarithmically at infinity, then $e^{-m}$ clearly belongs to $L^{p}$ for sufficiently large $p$. The consequences of this are summarized in the following result, where we assume that $m(\xi)\geq c\log(|\xi|^{2})$ for $|\xi|$ large; we may take $c=1$ without loss in generality.
@@
  • If $1\leq p<2$, satisfies that $u(t)\in L^{\frac{(N+\varepsilon)p}{(N+\varepsilon)-2pt}}(\mathbb{R}^{N})$ for $t\geq\frac{(N+\varepsilon)}{4}$. In particular, $u(t)\in L^{\infty}(\mathbb{R}^{N})$ for $t>\frac{N}{2p}$.
@@
$\psi_{1}^{\ell}(r)=\int_{r}^{1}\frac{\ell(s)}{s}\,ds,\qquad\psi_{2}^{\ell}(r)=\frac{1}{r^{2}}\int_{0}^{r}s\ell(s)\,ds.$
@@
then we have the lower estimate
@@
and some $c_{3}>0$ depending on $N$ and $\widetilde{J}$; see [33], also [11]. In the particular case $\ell\geq\mu$ near the origin for some positive constant, i.e., if $\widetilde{J}(z)\geq\mu|z|^{-N}$ for $|z|<1$, then
$m(\xi)\geq\mu\int_{|y|<1}\frac{1-\cos(y\cdot\xi)}{|y|^{N}}\,dy=\mu\int_{|z|<|\xi|}\frac{1-\cos(z_{1})}{|z|^{N}}\,dz=\mu\left(\frac{2\pi^{N/2}}{\Gamma(N/2)}\log|\xi|+O(1)\right),$
@@
Let $\mathcal{L}$ be a translation-invariant Lévy operator. If $\widetilde{J}(z)\geq\mu|z|^{-N}$ for $|z|<1$, the solution $u$ of problem ($\textup{P}_{\mathcal{L}}$) satisfies the conclusions of Section 6 with times depending on $\mu$ and $N$.
The previous results relied on studying the fundamental solution through its Fourier transform, and they have two clear drawbacks, both apparent in Section 6:
@@
thanks to the Picard-Lindelöf theorem. Since $\mathcal{W}_{in}$ is a maximal monotone operator, it is the infinitesimal generator of a solution semigroup $T(t)$, see the Appendix. Let $v(t)=T(t)w(t)$, and note that it satisfies the equation
@@
with $v(\cdot,0)=u_{0}$, using that translation-invariant operators commute. Finally, define $u(t)=(I-\Delta)^{-t}v(t)$. Since $v(t)\in L^{2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N})$ for all $t>0$, then $u(t)\in L^{2}_{2t}(\mathbb{R}^{N})\cap L^{p}_{2t}(\mathbb{R}^{N})$. On the other hand, using again that translation-invariant operators commute, we find that $u$ satisfies problem ($\textup{P}_{\mathcal{L}}$) with initial data $u_{0}$. Then, using that $\mathcal{L}$ is self-adjoint, we can use a standard approximation argument to prove that we can just take $u_{0}\in L^{2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N})$; see [12, Theorem 7.7].
@@
Thanks to the previous result and the embeddings used in Section 4.3, we can obtain the optimal time threshold after which solutions get bounded, which is $t>N/2p$.
With all the tools developed thus far, in this translation-invariant setting we can answer the question “is $\ell(r)\sim c$ for $r\sim 0$ the threshold for eventual ultracontractivity?”, and the answer is yes. Let us summarize all of this in the result below.

### v2

@@ 1.4. Outline of the paper
In Section 2 we introduce several necessary concepts and recall some useful tools already available in the literature. Section 3 is devoted to proving the equivalence between smoothing effects and families of logarithmic Sobolev inequalities. In Section 4 we focus on the model operator $\log(I-\Delta)$, applying the equivalence established above to obtain a new family of logarithmic Sobolev inequalities. Section 5 addresses general operators, making extensive use of the results from the preceding sections. In Section 6 we establish eventual ultracontractivity under the additional assumption of translation invariance, and we identify the borderline behaviour $|z|^{-N}$ at the origin as the threshold for this gradual regularization phenomenon. Finally, Section 7 contains additional comments and possible extensions. For completeness, an Appendix summarizes the relevant existence theory.
@@ 2.2. Concept of solution
2.3. Weak Lebesgue spaces and weak Young’s inequality
We define the weak Lebesgue space $L^{p,\infty}(\mathbb{R}^{N})$ (also known as the Marcinkiewicz space) as the set of measurable functions $f$ defined in $\mathbb{R}^{N}$ such that
(2.8) $\|f\|_{p,\infty}=\sup\left(\lambda\mu\{x\in\mathbb{R}^{N}:|f(x)|>\lambda\}^{1/p}\right)<\infty,$
where $\mu(A)$ denotes the measure of the set $A$. The functional $\|\cdot\|_{p,\infty}$ defines a quasinorm in this space. Weak Lebesgue spaces are heavily used in this work combined with weak Young’s inequality, which we state below.
The usual Young’s inequality states that, for $p,q,r\geq 1$ satisfying
(2.9) $1+\frac{1}{r}=\frac{1}{p}+\frac{1}{q},$
and $f\in L^{p}(\mathbb{R}^{N})$, $g\in L^{q}(\mathbb{R}^{N})$, the convolution fgf*g verifies
(2.10) $\|f*g\|_{r}\leq\|f\|_{p}\|g\|_{q}.$
If $p,q,r>1$, this inequality can be extended to weak Young’s inequality
(2.11) $\|f*g\|_{r}\leq c\|f\|_{p,\infty}\|g\|_{q}$
valid for all $f\in L^{p,\infty}(\mathbb{R}^{N})$, $g\in L^{q}(\mathbb{R}^{N})$, where $c>0$ is a constant depending on the exponents and the dimension. A natural question arises in this context: is weak Young’s inequality optimal? That is, assume that for any $r>q>1$ there exists some $C>0$ such that
(2.12) $\|f*g\|_{r}\leq C\|g\|_{q}$
for any $g\in L^{q}(\mathbb{R}^{N})$. Does this imply that $f$ belongs to $L^{p,\infty}(\mathbb{R}^{N})$ for the corresponding $p$ satisfying (2.9)? The result trivially holds under the assumption that $r=\infty$, see Section 6. Then, we will prove in Section 6 that this statement also holds if $f$ is a nonnegative radial function, nonincreasing in $|x|$. We believe that this result is likely available in the literature, although we have been unable to find a reference. We consider this lemma to be of independent interest.
@@ 3. Hypercontractivity and logarithmic Sobolev inequalities
  • pp–hypercontractive if there exist functions $q:[0,t_{\infty})\to[p,\infty)$ and $A_{p}:[0,t_{\infty})\to[0,\infty)$ for some $t_{\infty}\in(0,\infty]$ and $p\in[1,\infty)$, such that
    (3.1) $\|u(t)\|_{q(t)}\leq A_{p}(t)\|u_{0}\|_{p}\quad\text{for }0\leq t<t_{\infty},$
@@
There is a reason to talk about pp–hypercontractivity and not about pp–supercontractivity or pp–ultracontractivity. In the usual instances found in the literature, two pp–hypercontractivities with two different values of $p$ imply each other quantitatively, since they are both equivalent to the standard 22–logarithmic Sobolev inequality. However, to show strong hypercontractivity we need a family of independent pp–logarithmic Sobolev inequalities; therefore, the pp–hypercontractivities are also expected to be quantitatively independent; see the discussion below. The case of supercontractivity and ultracontractivity is different, since through an interpolation and iteration argument in (3.2) and (3.3), standard in the literature, one could substitute the 22–norm by any pp–norm.
@@
Let now $u$ be a solution to ($\textup{P}_{\mathcal{L}}$). Recall that we may assume that $u\geq 0$, and we will do so in the sequel. Also, all the integrals are understood to be over $\mathbb{R}^{N}$. Differentiating $\int u^{q(t)}(t)$, we formally obtain
@@
An approximation argument using continuous piecewise linear functions of $s$ and Section 3 shows that (3.22) is also true in $L^{1}$. Hence, (3.21) holds in $L^{1}$. ∎
@@
Now, let us prove a reciprocal result, which requires stronger hypotheses. While we required pp–hypercontractivity to hold in order to prove the pp–logarithmic Sobolev inequality, we need all the rr–logarithmic Sobolev inequalities for $r\geq p$ to prove the pp–hypercontractivity.
@@
Fix $p\geq 2$. Assume that the rr–logarithmic Sobolev inequalities (3.10) hold for all $r\geq p$. Then $\mathcal{L}$ is pp–hypercontractive with the family of exponents satisfying
@@
(3.37) $U^{\prime}(t)+s(t)\mathcal{E}\big(u^{s(t)-1}(t),u(t)\big)\leq\frac{s^{\prime}(t)}{s(t)}\left\{C(s(t))U(t)+D(s(t))\mathcal{E}\big(u^{s(t)-1}(t),u(t)\big)+U(t)\log\left(U(t)\right)\right\}.$
@@
Notice that, given $p\geq 2$, satisfying the rr–logarithmic Sobolev inequalities for every $r\geq p$ is equivalent to being rr–hypercontractive for every $r\geq p$. We would like to extend the result to $p\in(1,2)$. This will be done through a duality argument, which requires showing that the semigroup $e^{-t\mathcal{L}}$ is self-adjoint, an immediate consequence of $\mathcal{L}$ being self-adjoint.
@@
Let $u$ and $v$ be solutions to problem ($\textup{P}_{\mathcal{L}}$) with initial data $u_{0},v_{0}\in L^{2}(\mathbb{R}^{N})$ respectively. We wish to prove that
@@
Hence, we have that $t_{q^{\prime}}^{p^{\prime}}=t_{p}^{q}$, and we can define $A_{q^{\prime}}(t)=A_{p}(t)$. Therefore, $\mathcal{L}$ is $p$-hypercontractive for every $1<p\leq 2$.
@@
The function $q$ solution of the ODE satisfies the conditions of Theorem 3.11. Condition (3.51) is nothing but the standard blow-up condition for ODEs (also known as Osgood’s condition) written in terms of $D$ instead of $h$. ∎
@@
Under the hypotheses of Theorem 3.12, the operator $\mathcal{L}$ is eventually ultracontractive if, for any $p\geq 2$,
@@
(3.56) $\|u(t)\|_{q(t)}\leq A_{p}(t)\|u_{0}\|_{p},\quad\text{where }A_{p}(t)=\frac{q(t)}{p}.$
Hence, $\lim_{t\to t_{\infty}}A_{p}(t)=\infty$. This means that even though $q(t)$ blows up at $t_{\infty}$, we cannot deduce from this that $u$ becomes bounded. In fact, along the paper we will find that solutions in general will not become bounded at $t=t_{\infty}$. Nevertheless, they will become bounded for $t>t_{\infty}$, at least in the translation-invariant setting.
@@ 4. The $\log(I-\Delta)$ operator
To apply the results of the previous section, we will work with $0^{+}$-order operators. The idea is to work with an operator $\mathcal{L}$ with symbol $m(\xi)\sim\log(\xi)$ for $\xi\sim\infty$, since these are the kind of operators expected to have hypercontractivity, but not ultracontractivity; see Section 6. If we use $m(\xi)=\log(1+|\xi|^{2})$, we find the model operator for this work, $\mathcal{L}=\log(I-\Delta)$. This operator has the advantage of being a subordinated operator to the Laplacian. It is defined as
@@
It also appears as the infinitesimal generator of the Bessel potential $(I-\Delta)^{-t}$ semigroup, i.e.,
(4.2) $\log(I-\Delta)=\lim_{t\to 0^{+}}\frac{(I-\Delta)^{-t}-I}{t},$
which will be of importance in order to understand the gradual improvement of solutions.
@@ 4.1. Representation in terms of a Lévy kernel
On the other hand, using Laplace’s method (which estimates integrals via a Taylor’s expansion of the integrand; see for instance [43]), we can calculate, for large $|z|$,
$\begin{array}[]{rll}J(z)&=\displaystyle(4\pi)^{-\frac{N}{2}}\int_{0}^{\infty}t^{-\frac{N}{2}-1}e^{-\frac{|z|^{2}}{4t}-t}\,dt=\int_{0}^{\infty}g_{1}(t)e^{-g_{2}(z,t)}\,dt\\[8.53581pt] &\displaystyle=g_{1}(t_{0})e^{-g_{2}(z,t_{0})}\left(\int_{0}^{\infty}e^{-\frac{1}{2}\frac{\partial^{2}g_{2}}{\partial t^{2}}(z,t_{0})s^{2}}\,ds+o(1)\right)\\[8.53581pt] &\displaystyle=g_{1}(t_{0})e^{-g_{2}(z,t_{0})}\left(\left(\frac{2\pi}{\frac{\partial^{2}g_{2}}{\partial t^{2}}(z,t_{0})}\right)^{1/2}+o(1)\right),\end{array}$
@@
$|z|^{\frac{N+1}{2}}e^{|z|}J(z)=\left((2\pi)^{-\frac{N-1}{2}}+o(1)\right)\quad\text{as }|z|\to\infty.\qed$
The behaviour $J(z)\sim|z|^{-N}$ for $|z|\sim 0$ shows that this is a $0^{+}$-order operator. This fact is of great importance, since $\log(I-\Delta)$ will be the model for all operators whose kernel behaves that way. The behaviour at infinity is also studied in [55]. In particular the integral is a representation of a Bessel function of second order.
@@ 4.3. Smoothing properties
From this it is easy to observe that $H_{t}\in L^{\frac{N}{N-2t},\infty}(\mathbb{R}^{N})$, since $\||x|^{-N+2t}\|_{\frac{N}{N-2t},\infty}=\left(\frac{\omega_{N-1}}{N}\right)^{\frac{N-2t}{N}}$, where $\omega_{N-1}$ is the measure of the unit sphere in $\mathbb{R}^{N}$. With this information we may deduce the Sobolev embeddings found in Section 4.3. We may also deduce a similar result for $L^{1}$ initial data.
@@
Even though (4.21) was useful for this qualitative result, there are hidden factors depending on $t$ in its last term. Thus, the above formula is not enough to get precise quantitative estimates. Instead, we use a simpler but, for our purposes, much more powerful approach. In (4.14) we perform the change of variables $\tau=\frac{|x|^{2}}{4s}$ , which yields
@@ 4.4. Quantitative hypercontractivity
(4.36) $\|u(t)\|_{\frac{Np}{N-2pt}}\leq A_{p}(t)\|u_{0}\|_{p}\quad\text{for }0\leq t<\frac{N}{2p},$
@@
where $C_{N,p}(t)$ is the best possible constant for the inequality. We need the explicit expression of $C_{N,p}(t)$, or at least a sufficiently good upper bound. As they explain in [40], the best constant in weak Young’s inequality is exactly the best constant $c_{N,p,\lambda}$ for Hardy-Littlewood-Sobolev’s inequality,
@@ 4.5. Logarithmic Sobolev inequalities
We have now all the ingredients needed to establish the logarithmic Sobolev inequalities for $\log(I-\Delta)$. Let us denote by $H_{\mathcal{\log}}(\mathbb{R}^{N})$ the energy space associated to $\log(I-\Delta)$.
@@
(4.44) $A_{2}^{\prime}(t)=A_{2}(t)\Big[-\psi\!\left(\frac{N}{2}-2t\right)-\psi\!\left(\frac{N}{2}+2t\right)-\ln(4\pi)+\frac{2}{N}\ln\!\left(\frac{\Gamma(N)}{\Gamma(\tfrac{N}{2})}\right)\Big],$
@@ 4.6. Comparison with the operator $\log(-\Delta)$
with the same function coefficient $A_{p}$ obtained in Theorem 4.7. Hence, one may try to apply Theorem 3.11 in order to obtain the family of logarithmic Sobolev inequalities for $\log(-\Delta)$. However, the expression (3.15) may not be well-defined for $u(t)=R_{t}*u_{0}$, since it does not belong to the appropriate spaces. Consequently, obtaining logarithmic Sobolev inequalities —at least by our method— is not possible. In fact, $u(t)$ cannot be considered a solution in the sense of Section 2.2, since $u(t)$ may not belong to $L^{2}$ for $t>0$. It is clear that Hille-Yosida’s Theorem, Theorem 7.1, cannot be applied here, since $\log(-\Delta)$ is not a monotone operator.
Even more important, solutions $u(t)=R_{t}*u_{0}$ for $u_{0}\in L^{p}(\mathbb{R}^{N})$ cease to be well defined for $t>\frac{N}{2p}$, even as tempered distributions, since the Fourier multiplier is not locally integrable anymore. So the main differences between problem ($\textup{P}_{\mathcal{L}}$) for $\log(-\Delta)$ and $\log(I-\Delta)$ are:
@@ 6. Translation-invariant operators
Let $1\leq p<\infty$. If $e^{-m}\in L^{p}(\mathbb{R}^{N})$, then $\mathcal{H}_{t}\in L^{1}(\mathbb{R}^{N})\cap L^{\frac{p}{p-t}}(\mathbb{R}^{N})$ for $t\in[p/2,p)$. In particular, $\mathcal{H}_{t}\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ for $t\geq p$.
@@
Define $\tau_{z}f(x)=f(z-x)$. Then, condition (6.5) implies that
(6.6) $|\langle H_{t_{0}},\tau_{0}f\rangle|\leq C\|\tau_{0}f\|_{p}$
for every $f\in L^{p}(\mathbb{R}^{N})$, since $\tau_{0}$ is an isometry in $L^{p}(\mathbb{R}^{N})$. Hence, renaming $g=\tau_{0}f$, and taking the supremum of $g\in L^{p}(\mathbb{R}^{N})$ such that $\|g\|_{p}=1$, we conclude what we wanted. ∎
If $m$ grows logarithmically at infinity, then $e^{-m}$ clearly belongs to $L^{p}$ for sufficiently large $p$. The consequences of this are summarized in the following result, where we assume that $m(\xi)\geq c\log(|\xi|^{2})$ for $|\xi|$ large; we may take $c=1$ without loss of generality.
@@
  • If $1\leq p<2$, $u(t)\in L^{\frac{(N+\varepsilon)p}{(N+\varepsilon)-2pt}}(\mathbb{R}^{N})$ for $\frac{(N+\varepsilon)}{4}\leq t\leq\frac{N}{2}$. In particular, $u(t)\in L^{\infty}(\mathbb{R}^{N})$ for $t>\frac{N}{2p}$.
@@
$\ignorespaces\ignorespaces\psi_{1}^{\ell}(r)=\int_{r}^{1}\frac{\ell(s)}{s}\,ds,\qquad\psi_{2}^{\ell}(r)=\frac{1}{r^{2}}\int_{0}^{r}s\ell(s)\,ds.$
@@
we have the lower estimate
@@
for some $c_{3}>0$ depending on $N$ and $\widetilde{J}$; see [37], and also [11]. In the particular case in which $\ell\geq\nu$ near the origin for some positive constant, i.e., if $\widetilde{J}(z)\geq\nu|z|^{-N}$ for $|z|<1$, then
$m(\xi)\geq\nu\int_{|y|<1}\frac{1-\cos(y\cdot\xi)}{|y|^{N}}\,dy=\nu\int_{|z|<|\xi|}\frac{1-\cos(z_{1})}{|z|^{N}}\,dz=\nu\left(\omega_{N-1}\log|\xi|+O(1)\right),$
@@
Let $\mathcal{L}$ be a translation-invariant Lévy operator. If $\widetilde{J}(z)\geq\nu|z|^{-N}$ for $|z|<1$ and $\nu>0$, the solution $u$ of problem ($\textup{P}_{\mathcal{L}}$) satisfies the conclusions of Section 6 with times depending on $\nu$ and $N$.
The previous results relied on studying the fundamental solution through its Fourier transform, and they have two clear drawbacks, already apparent in Section 6:
@@
whose existence is guaranteed by Picard-Lindelöf’s theorem. Since $\mathcal{W}_{in}$ is a maximal monotone operator, it is the infinitesimal generator of a solution semigroup $T(t)$, see the Appendix. Let $v(t)=T(t)w(t)$, and note that it satisfies the equation
@@
with $v(\cdot,0)=u_{0}$ (this is proved using that translation-invariant operators commute). Finally, define $u(t)=(I-\Delta)^{-t}v(t)$. Since $v(t)\in L^{2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N})$ for all $t>0$, then $u(t)\in L^{2}_{2t}(\mathbb{R}^{N})\cap L^{p}_{2t}(\mathbb{R}^{N})$. On the other hand, using again that translation-invariant operators commute, we find that $u$ satisfies problem ($\textup{P}_{\mathcal{L}}$) with initial data $u_{0}$. It is straightforward to check that it is the unique solution: just use the theory of Theorem 7.1 with $\mathcal{L}=\mathcal{W}^{\textup{in}}$, adding a right-hand side $\mathcal{W}^{\textup{out}}u$, which can be managed be means of Gronwall’s inequality. Then, since $\mathcal{L}$ is self-adjoint, we can use a standard approximation argument to prove that we can take $u_{0}\in L^{2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N})$; see [12, Theorem 7.7].
@@
Thanks to the previous result and the embeddings used in Section 4.3, we can obtain the optimal time threshold after which solutions get bounded, which is $t>N/2p$. Observe that the result is valid for anisotropic kernels $\widetilde{J}$.
Finally, we address the question of whether $\widetilde{J}(z)\sim|z|^{-N}$ for $|z|\sim 0$ constitutes the threshold for the gradual smoothing phenomenon. The answer is affirmative: any radial perturbation that increases the singularity of the kernel renders the operator ultracontractive, whereas any radial perturbation that decreases it prevents any improvement of the fundamental solution.